Niagara falls are 50m high. calculate the potential energy of 0.1 cubic meter of water at the top relative to the bottom. density of water is 1000 kg m^-3. take g= 10m/s^2
1
Expert's answer
2019-05-31T09:44:06-0400

Potential energy is:

"E=mgh"

Mass:

"m=\rho*V"

then:

"E=\rho Vgh=1000\times 0.1\times 10 \times 50 = 50000 J = 50 kJ"